How can I solve this? At December 31, 2016, House Co. reported the following information on its balance sheet. Accounts receivable $960,000 Less: Allowance for doubtful accounts 80,000 During 2017, the company had the following transactions related to receivables. 1. Sales on account $3,700,000 2. Sales returns and allowances 50,000 3. Collections of accounts receivable 2,810,000 4. Write-offs of accounts receivable deemed uncollectible 90,000 5. Recovery of bad debts previously written off as uncollectible 29,000 Instructions (a) Prepare the journal entries to record each of these fi ve transactions. Assume that no cash discounts were taken on the collections of accounts receivable. (b) Enter the January 1, 2017, balances in Accounts Receivable and Allowance for Doubtful Accounts, post the entries to the two accounts (use T-accounts), and determine the balances. (c) Prepare the journal entry to record bad debt expense for 2017, assuming that an aging of accounts receivable indicates that expected bad debts are $115,000. (d) Compute the accounts receivable turnover for 2017 assuming the expected bad debt information provided in (c).
Answer #1

Answer (a):

Answer b:

Answer c:

Allowance for doubtful accounts required = $115,000

Unadjusted Credit balance in Allowance for doubtful accounts = $19,000

Bad debt expense = 115000 - 19000 =$96,000

Answer d:

Dec 31, 2016 Net receivable = $960,000 - 80,000 =$880,000

Dec 31, 2017 Net receivable = $1,710,000 - 115,000 =$1,595,000

Average receivable = $1,237,500

Net Sales = $3,700,000 - 50000 = $3,650,000

Accounts receivable turnover for 2017 = Net Sales on account / Average receivable = 3,650,000 / 1,237,500 = 2.95

Accounts receivable turnover for 2017 = 2.95